What is the kinetic energy of a thrown 13.2 N stick moving at 3.89 m/s?


1
Expert's answer
2021-03-08T08:24:01-0500

13.2 N = 1.34 kg

Kinetic energy "E_k = \frac{1}{2}mv^2"

"= \frac{1}{2} \times 1.34 \times (3.89)^2 \\

= 10.13 \;J"

Answer: 10.13 J